# Shelf-Supervised Multi-Modal Pre-Training for 3D Object Detection

### v1

@@ Appendix B Ablation on Self-Training
Self-Training Algorithm. Given Kannotated training data and (1-K%) unlabeled training data with CM3D pseudo-labels, we first pre-train a randomly initialized detector on (1-K%) pseudo-labels, fine-tune on Kannotated data, and use the resulting fine-tuned model to re-label the (1-K%) unlabeled training data. We iterative refine the (1-K%) pseudo-labels through multiple rounds of self-training. Importantly, we randomly initialize the detector after each round of self-training and pseudo-label generation to simplify training. We find that even one round of self-training significantly improves NDS when fine-tuning detectors with limited annotations. Additional rounds of self-training provide limited improvements.
@@
Appendix C Ablation on Copy-Paste Augmentation
@@
Appendix D nuScenes Class Agnostic Evaluation
@@
Appendix E Replacing Detic with GroundingDINO
@@
Appendix F Comparing ChatGPT Priors vs. Real Shape Priors
@@
Appendix G More Qualitative Results
@@
Appendix H Evaluating CM3D on Waymo
@@
Appendix I Limitations and Future Work

### v2

@@ Appendix B Ablation on Self-Training
Figure 6: Self-Training Procedure. We train a detector with ground truth annotations from Kof the train-set an pseudo-labels from (1-K%) of the train-set. We iteratively use the trained detector to predict better pseudo-labels on the (1-K%) subset until the accuracy of the predicted detections plateaus. We visualize one round of self-training above.
@@
Self-Training Algorithm. As shown in Fig 6, given Kannotated training data and (1-K%) unlabeled training data with CM3D pseudo-labels, we first pre-train a randomly initialized detector on (1-K%) pseudo-labels, fine-tune on Kannotated data, and use the resulting fine-tuned model to re-label the (1-K%) unlabeled training data. We iterative refine the (1-K%) pseudo-labels through multiple rounds of self-training. Importantly, we randomly initialize the detector after each round of self-training and pseudo-label generation to simplify training. We find that even one round of self-training significantly improves NDS when fine-tuning detectors with limited annotations. Additional rounds of self-training provide limited improvements.
@@
Appendix C Ablation on Scaling Up Pseudo-Labels
We demonstrate how our method scales with an increasing amount of unlabeled data (shown as a percentage of the full nuScenes train-set) with a fixed budget of 5% labeled data in Fig. 7. For simplicity, we do not perform self-training.
Figure 7: Scaling Up Pseudo-Labels. We plot the impact of scaling up the amount of unlabeled data on the nuScenes dataset on semi-supervised detection performance (mAP). Increasing the number of pseudo-labels (PL, shown as a percentage of the nuScenes train-set) used for pre-training significantly improves semi-supervised detection accuracy. This suggests that our method will continue to improve with larger sets of unlabeled datasets.
Appendix D Ablation on Copy-Paste Augmentation
@@
Appendix E nuScenes Class Agnostic Evaluation
@@
Appendix F Replacing Detic with GroundingDINO
@@
Appendix G Comparing ChatGPT Priors vs. Real Shape Priors
@@
Appendix H More Qualitative Results
@@
Appendix I Evaluating CM3D on Waymo
@@
Appendix J Limitations and Future Work
@@
Future Work: Application to Broader Robotics Tasks. Our approach leverages domain-specific insights, and may not be directly generalizable to robotics applications without LiDAR or HD Maps. Future work can adapt our pseudo-labeling framework for such situations. For example, instead of using LiDAR, one can directly use metric depth estimates [piccinelli2024unidepth] from state-of-the-art monocular depth estimators. Similarly, one can estimate orientation with rotating calipers or track-level information. We present preliminary results of our pseudo-labeler applied to non-AV applications in Fig. 9.
Figure 9: Generalized Pseudo-Labeler. We visualize the output of our generalized psuedo-labeler on several images above. Importantly, we don’t assume we have access to LiDAR for precise depth estimates or HD maps for orientation estimation. Instead, we use UniDepth [piccinelli2024unidepth] to generate a pseudo-point cloud and estimate orientation using rotating calipers. Importantly, orientation estimates can be up to 180 degrees off because this current setup cannot estimate heading. This suggests that our pseudo-labeler can generalize beyond autonomous vehicle settings with minor modifications. We expect that semi-supervised training can further improve the quality of predicted bounding boxes.